OpenStudy (anonymous):

Solve the following system of equations: x - 2y = 14 x + 3y = 9 A) (1, 12) B) (-1, -12) C) (12, -1) D) (12, 1)

pooja195 (pooja195):

We want to cancel out each varible to find the value x - 2y = 14 x + 3y = 9 A negative - a positvive results in a 0 so whats the opposite of 1 ?

OpenStudy (anonymous):

-1

pooja195 (pooja195):

Good -1(x - 2y = 14) x + 3y = 9 \[\huge~\rm~-1(x - 2y = 14) \] distribute

OpenStudy (anonymous):

how would I do that

pooja195 (pooja195):

-1 times x = ? -1 times -2 =? -1 times 14=?

OpenStudy (anonymous):

i don't know...

pooja195 (pooja195):

If you dont want to work im not helping you i did 1 with you and nearly gave u the answer if you dont want to work and are looking for JUST answers this is the wrong site.